Find Io using mesh analysis. 10 12/0° v 2/0°A 31.2 + T-jin
Answer #1

Given 2 22oA - 12-24 Io-= 2 1220V x 12 dr. in a I Il 1 T-12 lo a > ☺ outer roup Let the current flowing are I2, I, to from thi substituting I, Ig in equation ② 60+3)-. GMP) +-2 +1o = 0 Io (1+1 6 + ) = -6(1+3)+2 - --6-6j +2 To (E ) 5-65-4 To = -4 (33